A rock is thrown upwards with a speed of 7m/s. How much time does it take to reach the peak of its arc


1
Expert's answer
2022-05-24T13:56:14-0400
"t=\frac{v_0}{g}\\
=\rm\frac{7\:m/s}{9.8\: m/s^2}=0.71\:s"